# Bipartite Fluctuations of Critical Fermi Surfaces

### v2

@@ Appendix E Dual Theories for CFL-FL Transition
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
$|\textrm{D}_{\tilde{b}}\varphi_{1}|^{2}+|\varphi_{1}|^{4}+|\textrm{D}_{\tilde{c}}\varphi_{2}|^{2}+|\varphi_{2}|^{4}-\frac{\mathtt{i}}{4\pi}\tilde{b}\textrm{d}\tilde{b}-\frac{\mathtt{i}}{4\pi}\tilde{c}\textrm{d}\tilde{c}-\frac{\mathtt{i}}{2\pi}\tilde{a}\textrm{d}(a+\tilde{b}+\tilde{c}).$ (110)
@@
Appendix F Static Structure Factor at Criticality
@@
In this appendix, we provide technical details regarding the evaluation of integrals for the static structure factors, as expressed by Eq. (37) and Eq. (41), at the quantum critical points. If we naively consider the small-$\bm{k}$ expansion of the integrand before doing the integral, the leading-order term can integrated analytically, resulting in
@@
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})\approx\int\frac{\textrm{d}\omega}{2\pi}\frac{-|\bm{k}|^{2}k_{F}v_{F}(k_{F}v_{F}\sigma_{xx}^{b}+4\pi\omega(\sigma_{xx}^{b})^{2}-4\pi\omega(\sigma_{xy}^{b})^{2})}{\omega(k_{F}v_{F}+4\pi\omega\sigma_{xx}^{b}){}^{2}-16\pi^{2}\omega^{3}(\sigma_{xy}^{b})^{2}}=-\frac{\sigma_{xx}^{b}}{\pi}|\bm{k}|^{2}\log(1/|\bm{k}|)+\ldots.$ (112)
@@
Here, a logarithmic divergence arises from the $\omega$-integral, leading to the factor $\log(1/|\bm{k}|)$ due to the identical scaling of $\omega$ and $\bm{k}$ at the critical point. The naive analytical calculation of Eq. (41) proceeds in a similar manner and is a special case of Eq. (112) with $\sigma_{xy}=0$.
@@
Figure 4: (1) The numerical result (in black) of the static structure factor given by Eq. (37) at the CFL-FL Transition, where $k_{F}=\sqrt{2\pi}$, $v_{F}=0.25$, and $\sigma_{xx}=-\sigma_{xy}=1/(8\pi)$; (2) The numerical result (in black) of the static structure factor given by Eq. (41) at the Mott Transition, where $k_{F}=\sqrt{2\pi}$, $v_{F}=0.25$, and $\sigma_{xx}=0.355155/(2\pi)$. The numerical results are compared to the CFT-like scaling (in red) and that of ordinary Fermi surfaces (in blue).
@@
To confirm the final result in Eq. (112), we also numerically evaluated the integrals using the exact expressions from Eq. (37) and Eq. (41), with the results presented in Fig. 4. In these numerical calculations, the value of $k_{F}$ is determined by the Luttinger theorem $V_{\textrm{FS}}/(2\pi)^{2}=1/2$ at half-filling [^6] The Luttinger theorem, as it applies to the composite Fermi liquid and the spinon Fermi surface, is discussed in Ref. [45]. The more precise expressions are $V_{\textrm{FS}}/(2\pi)^{2}=(1/2)/(2\pi\ell_{B}^{2})$ and $V_{\textrm{FS}}/(2\pi)^{2}=(1/2)/V_{\textrm{cell}}$, where $\ell_{B}$ represents the magnetic length and $V_{\textrm{cell}}$ denotes the volume of the unit cell. Our purpose is to numerically verify the analytic result in Eq. (112). To this end, we set $12\pi\ell_{B}^{2}=$ 1 and $V_{\textrm{cell}}=1$ for simplicity., where the Fermi-surface volume is given by $V_{\textrm{FS}}=\pi k_{F}^{2}$. We use a small value of $v_{F}$ for estimations. For the chargon conductivity $\sigma_{xx}^{b}$ at the Mott transition, we adopt the critical conductivity at the XY transition from conformal bootstrap [101]. For chargons at the CFL-FL transition, as a very crude estimation, we assume the universal transverse and Hall resistivities $\rho_{xx}^{b},\rho_{xy}^{b}$ are both around $4\pi$ [56]. In both cases, the numerical results agree well with the analytical expressions at small-$\bm{k}$, confirming the expectation that the static structure factor of critical Fermi surfaces behaves in a CFT-like manner and deviates from that of ordinary Fermi surfaces.
Appendix G Non-Linear Bosonization
@@
[…]

### v3

@@ Appendix F Dual Theories for CFL-FL Transition
$\bar{\psi}\not{\textrm{D}}_{A}\psi\quad\longleftrightarrow\quad\bar{\chi}\not{\textrm{D}}_{a}\chi-\frac{\mathtt{i}}{2\pi}a\textrm{d}b+\frac{\mathtt{i}2}{4\pi}b\textrm{d}b-\frac{\mathtt{i}}{2\pi}b\textrm{d}A+\frac{\mathtt{i}}{4\pi}A\textrm{d}A,$ (118)
@@
$\bar{\psi}\not{\textrm{D}}_{A}\psi-\frac{1}{2}\frac{\mathtt{i}}{4\pi}A\textrm{d}A\quad\longleftrightarrow\quad\bar{\chi}\not{\textrm{D}}_{a}\chi-\frac{1}{2}\frac{\mathtt{i}}{4\pi}a\textrm{d}a-\frac{1}{2}\frac{\mathtt{i}}{2\pi}a\textrm{d}A.$ (119)
@@
$\bar{\psi}\not{\textrm{D}}_{A}\psi-\frac{1}{2}\frac{\mathtt{i}}{4\pi}A\textrm{d}A\quad\longleftrightarrow\quad\bar{\chi}\not{\textrm{D}}_{a}\chi-\frac{1}{2}\frac{\mathtt{i}}{4\pi}a\textrm{d}a+\frac{1}{2}\frac{\mathtt{i}}{2\pi}a\textrm{d}A.$ (120)
@@
$\sum_{I=1}^{2}\bar{\psi}_{I}\not{\textrm{D}}_{\tilde{a}}\psi_{I}+\frac{\mathtt{i}}{2\pi}\tilde{a}\textrm{d}(A-a)\quad\longleftrightarrow\quad\sum_{I=1}^{2}\bar{\chi}_{I}\not{\textrm{D}}_{\tilde{b}}\chi_{I}-\frac{2\mathtt{i}}{4\pi}\tilde{b}\textrm{d}\tilde{b}-\frac{\mathtt{i}}{2\pi}\tilde{b}\textrm{d}(A-a)-\frac{\mathtt{i}}{4\pi}(A-a)\textrm{d}(A-a).$ (122)
@@
$\bar{\psi}\not{\textrm{D}}_{A}\psi\quad\longleftrightarrow\quad|\textrm{D}_{a}\varphi|^{2}+|\varphi|^{4}-\frac{\mathtt{i}}{4\pi}a\textrm{d}a-\frac{\mathtt{i}}{2\pi}a\textrm{d}A.$ (123)
Using Eq. (123), we can express the dual theory of $\bar{\psi}_{1}\not{\textrm{D}}_{\tilde{a}}\psi_{1}+\bar{\psi}_{2}\not{\textrm{D}}_{\tilde{a}}\psi_{2}-\frac{\mathtt{i}}{2\pi}\tilde{a}\textrm{d}a$ as follows
@@
Appendix G Static Structure Factor at Criticality
@@
In this appendix, we provide technical details on evaluating the integrals for the static structure factors at the quantum critical points, as given by Eq. (40) and Eq. (44). If one naively expands the integrand at small $\bm{k}$ before performing the integration, the leading-order term can be analytically integrated, yielding
@@
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})\approx\int_{0}^{+\infty}\frac{\textrm{d}\omega}{\pi}\frac{|\bm{k}|^{2}k_{F}v_{F}(k_{F}v_{F}\sigma_{xx}^{b}+4\pi\omega(\sigma_{xx}^{b})^{2}-4\pi\omega(\sigma_{xy}^{b})^{2})}{\omega(k_{F}v_{F}+4\pi\omega\sigma_{xx}^{b}){}^{2}-16\pi^{2}\omega^{3}(\sigma_{xy}^{b})^{2}}=\frac{\sigma_{xx}^{b}}{\pi}|\bm{k}|^{2}\log(1/|\bm{k}|)+\ldots$ (126)
@@
where the logarithmic divergence arises from the $\omega$-integral. The factor $\log(1/|\bm{k}|)$ reflects the identical scaling of $\omega$ and $\bm{k}$ at the critical point. The analytic evaluation of Eq. (44) proceeds similarly and can be viewed as a special case of Eq. (126) with $\sigma_{xy}^{b}=0$.
@@
Figure 4: (1) Numerical result (black) for the static structure factor at the CFL-FL transition, based on Eq. (40), with parameters $k_{F}=\sqrt{2\pi}$, $v_{F}=0.25$, and $\sigma_{xx}^{b}=-\sigma_{xy}^{b}=1/(8\pi)$; (2) Numerical result (black) for the static structure factor at the Mott transition, based on Eq. (44), with $k_{F}=\sqrt{2\pi}$, $v_{F}=0.25$, and $\sigma_{xx}^{b}=0.355155/(2\pi)$. In both panels, the results are compared with the CFT-like scaling behavior (red) and that of ordinary Fermi surfaces (blue).
@@
To verify the analytic result in Eq. (126), we also numerically evaluated the integrals using the exact expressions in Eq. (40) and Eq. (44), with the results presented in Fig. 4. In these numerical calculations, the Fermi momentum $k_{F}$ is determined by the Luttinger theorem $V_{\textrm{FS}}/(2\pi)^{2}=1/2$ at half-filling [^8] The Luttinger theorem, as it applies to the composite Fermi liquid and the spinon Fermi surface, is discussed in Ref. [46]. The more precise expressions are $V_{\textrm{FS}}/(2\pi)^{2}=(1/2)/(2\pi\ell_{B}^{2})$ and $V_{\textrm{FS}}/(2\pi)^{2}=(1/2)/V_{\textrm{cell}}$, where $\ell_{B}$ represents the magnetic length and $V_{\textrm{cell}}$ denotes the volume of the unit cell. Our purpose is to numerically verify the analytic result in Eq. (126). To this end, we set $2\pi\ell_{B}^{2}=1$ and $V_{\textrm{cell}}=1$ for simplicity., where the Fermi-surface volume is given by $V_{\textrm{FS}}=\pi k_{F}^{2}$. We choose a small $v_{F}$ to facilitate the numerical estimation. For the chargon conductivity $\sigma_{xx}^{b}$ at the Mott transition, we adopt the critical conductivity value obtained from the conformal bootstrap for the 3D XY universality class [136]. For chargons at the CFL-FL transition, following Ref. [57], we crudely estimate that the universal transverse and Hall resistivities $\rho_{xx}^{b},\rho_{xy}^{b}$ are both of order $4\pi$. In both cases, the numerical results agree well with the analytic predictions at small $\bm{k}$, confirming that the static structure factor at critical Fermi surfaces exhibits CFT-like scaling and deviates from that of ordinary Fermi surfaces.
Appendix H Non-Linear Bosonization
@@
$F_{0}(\bm{x},\bm{k})\overset{\textrm{df}}{=}\langle\textrm{FS}|\mathbbm{t}(\bm{x},\bm{k})|\textrm{FS}\rangle=\Theta(\epsilon_{F}-\epsilon_{\bm{k}}),$ (129)